Object A has a mass of 40 kg and is traveling at 25 m/s [W] when it strikes a stationary object B with a mass of 29 kg in a perfectly elastic collision. Determine the velocity of object B after the collision.


1
Expert's answer
2021-03-23T11:10:39-0400

In a perfectly elastic collision the momentum conservation law holds. Thus, the total momentum before the collision:


"p = m_1v_1"

is equal to the total momentum after the collision:


"p' = m_1v_1' + m_2v_2'"

where "m_1 = 40kg, m_2 = 29kg" are the masses of the objects, "v_1 = 25m\/s" is the speed of the first object before the collision, and "v_1', v_2'" are their speeds after the collsion. Thus, obtain:


"m_1v_1 = m_1v_1' + m_2v_2'"


Futhermore, the energy conservation law also holds for an elastic collision:


"\dfrac{m_1v_1^2}{2} = \dfrac{m_1v_1'^2}{2} +\dfrac{m_2v_2'^2}{2}"

Expressing "v_1'" from the first equation and substituting it into the last one, after doing some algebra obtain:


"v_2'^2\left( \dfrac{m_2^2}{m_1} + m_2 \right)-2m_2v_1v_2'=0\\
v_2'^2\left( \dfrac{m_2}{m_1} +1 \right)-2v_1v_2'=0\\
v_2' \left( v_2'\left( \dfrac{m_2}{m_1} +1 \right) - 2v_1 \right) = 0"

One solution is "v_2' = 0", which is not possible. Another one is:

"v_2'\left( \dfrac{m_2}{m_1} +1 \right) - 2v_1 = 0\\
v_2' = \dfrac{2v_1}{\dfrac{m_2}{m_1} +1} = \dfrac{2m_1v_1}{m_1+m_2}"

Substituting the numbers, obtain:


"v_2'=\dfrac{2\cdot 40\cdot 25}{40+29} \approx 28.99 \space m/s"